(A) Oil flows in a pipe 100 mm bore diameter with a Reynolds number of 500. The density is 800 kg/m3. calculate the velocity of a streamline at a radius of 40 mm. the viscosity = 0.08 Ns/m2.

(B) Identify and discuss the features of external flow?
1
Expert's answer
2020-04-27T01:55:12-0400

(A) Reynold number of fluid flow=500 and its bore diameter=100mm, Density is = 800 "\frac{kg}{m^3}"

The coefficient of viscosity is 0.08 Ns/m2

we know that,

Re= "\frac{\rho vd}{\mu}"

"500= \frac{800 \times v \times 0.01}{0.08}"


v= 0.5 m/s

Now, for finding velocity at 40 mm dia we use Poiseulli equation for laminar flow

"u= \frac{ \Delta P (R^2-r^2)}{4L \mu}" , here "\Delta P= \frac{32 \mu L}{vd^2}=128 L"

"u=\frac { 128 L (0.05^2 -0.04^2)}{4L 0.08}"

u=0.36 m/s

this is the velocity of fluid at radius of 40 mm

(B) External flow is that type of flow which is highly viscous and develop to boundary layers it is due the the characteristic causing interaction between body and fluid.